OpenStudy (anonymous):

Solve (xy+y^2+x^2) dx -( x^2 )dy = 0 by substitution, any help?

OpenStudy (some.random.cool.kid):

sorry, not my style :{

ganeshie8 (ganeshie8):

it is homogeneous, so try \(\large \mathrm{y = vx}\)

OpenStudy (anonymous):

first, how do i verify if it is homogeneous? My book is very vague about that.

OpenStudy (anonymous):

is it if equal to 0?

ganeshie8 (ganeshie8):

replace `x` by `kx` and `y` by `ky`

ganeshie8 (ganeshie8):

(xy+y^2+x^2) dx -( x^2 )dy = 0 ((kx)(ky)+(ky)^2+(kx)^2) dx -( (kx)^2 )dy = 0 k^2 [(xy+y^2+x^2) dx -( x^2 )dy] = 0

ganeshie8 (ganeshie8):

Notce that entire k power got factored out, so we can use this for testing quickly whether a DE is homogenerous or not..

OpenStudy (anonymous):

ah I see. So then when I replace the whole thing by y=ux, I can then do (x^2u+u^2x^2+x^2)dx - x^2 dy and I can factor out x^2, and be left with u's only. right? also replacing dy = xdu + u dx right?

ganeshie8 (ganeshie8):

Exactly !

ganeshie8 (ganeshie8):

maybe it is easier to see this by putting it in `dy/dx = f(x,y)` form first

OpenStudy (anonymous):

hmmmm I don't know how to do that, or not sure rather*

ganeshie8 (ganeshie8):

\[(xy+y^2+x^2) dx -( x^2 )dy = 0\] is equivalent to \[\dfrac{dy}{dx} = \dfrac{xy+y^2+x^2}{x^2}\]

OpenStudy (anonymous):

Oh I see what you mean, my bad haha, stressing over an exam isn't helping to see clearly.

OpenStudy (anonymous):

but afterwards, I see that my solutions cancel out the x^2 from x^2(u+u^2+1)dx - x^2( x du + u dx) giving (u+u^2+1) dx - (x du + u dx). is it just because it is a comon coefficient between both functions?

OpenStudy (anonymous):

and then you can cancel everything and be left with u^2+1 dx - xdu = 0 and use separable equation..

OpenStudy (anonymous):

And lastly if I may ask, for substitution, how do you decide to use it? after verifying if it is not exact? and how did you know to use y = vx, just a rule type thing?

ganeshie8 (ganeshie8):

It helps to know that there are only two standard ways for solving ANY differential equation : 1) separable 2) linear We try to change all other types of equations into one of these known forms and solve..

ganeshie8 (ganeshie8):

look at your DE in this form : \[\dfrac{dy}{dx} = \dfrac{xy+y^2+x^2}{x^2}\]

OpenStudy (anonymous):

to me it doesn't seem linear with y^2

ganeshie8 (ganeshie8):

as you can see the right hand side is pretty complicated with a mix of x and y terms

ganeshie8 (ganeshie8):

but if you look closer, you will see that if you replace \(y\) by \(ux\), both \(x\) and \(y\) disappear giving you a rational expression in \(u\) alone : \[\dfrac{dy}{dx} = \dfrac{xy+y^2+x^2}{x^2}\] plugin \(y=ux \implies \frac{dy}{dx} = \frac{du}{dx}x+u \) the eqn becomes : \[\dfrac{du}{dx}x+u = \dfrac{x(ux)+(ux)^2+x^2}{x^2}\] \[\dfrac{du}{dx}x+u = \dfrac{x^2(u+u^2+1}{x^2}\] \[\dfrac{du}{dx}x+u = u+u^2+1\] which is separable

ganeshie8 (ganeshie8):

Here is a full solution

OpenStudy (anonymous):

soorry to bother you one last time but where did you get y=ux --> du/dx x+u and for the rest I think i could give you a hug for your patience and help, thank you for allowing me to understand this concept! :)

ganeshie8 (ganeshie8):

\(\large \color{magenta}{\heartsuit} \)\[y = ux\] differentiate both sides with.resepct.to \(x\)

ganeshie8 (ganeshie8):

use product rule. .